media/6e4/6e4c9b1e-adec-47d4-a5ce-e5
Answer #1
Concepts and reason

Nucleophilic substitution reactions involve the replacement of a weak nucleophile by a stronger one. This type of reactions can take place via two mechanisms. Biomolecular nucleophilic substitution reactions (SN2)\left( {{{\rm{S}}_{\rm{N}}}2} \right) and unimolecular nucleophilic substitution reactions (SN1)\left( {{{\rm{S}}_{\rm{N}}}1} \right) .

Fundamentals

Biomolecular nucleophilic substitution (SN2)\left( {{{\rm{S}}_{\rm{N}}}2} \right) is a second order reaction which involves only one step whereas, unimolecular nucleophilic substitution is a first order reaction having two steps.

Methoxy (CH3O)\left( {{\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{ - }}{{\rm{O}}^ - }} \right) group will attack the carbon atom attached to bromine because the electron density of methoxy group is very high. Therefore, it works as a nucleophile.

H
H3C—(
Br: +
Ö—CH —
H3C—
o—CH2 +
Br

Bromine leaves the carbon atom and replaced by methoxy group.

H3C-
Ć
—Br:
+
:0—
CHz -
Học
_¢_5_CH
+
Br

Ans:

Curved arrow representation of the reaction is as follows:

HOTO: +
H|3|H
HC-C
|